ganeshie8 (ganeshie8):

A nonuniform bar is suspended at rest in a horizontal position by two massless cords as shown. One cord makes the angle θ = 36.9° with the vertical; the other makes the angle \(\phi\) = 53.1° with the vertical. If the length L of the bar is 6.10 m, compute the distance x from the left end of the bar to its center of mass.

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

I'm not sure @Michele_Laino

Parth (parthkohli):

Net force = 0:\[T_1 \cos \theta + T_2 \cos \phi=mg\]\[T_1 \sin \theta = T_2 \sin \phi\]Net torque about any point is zero. I'm going to take that point as the center of mass.\[T_1 \cos \theta(x) = T_2 \cos \phi (L-x)\]

Parth (parthkohli):

And 37-53-90 are the angles of the 3-4-5 (or similar derivative) triangles.

ganeshie8 (ganeshie8):

so from first equations, i suppose, we have T1 and T2

Parth (parthkohli):

Yeah, and then we plug and play those values in the equation for torque.

ganeshie8 (ganeshie8):

Awesome! let me plugin the numbers... it feels nice to see the textbook answer match with mine

Parth (parthkohli):

Which textbook is this?

ganeshie8 (ganeshie8):

Hey, we don't know mass of the rod ?

ganeshie8 (ganeshie8):

Net force = 0:\[T_1 \cos \theta + T_2 \cos \phi=\color{red}{m}g\]\[T_1 \sin \theta = T_2 \sin \phi\]

Parth (parthkohli):

You're right... this is not as straightforward as I thought. If we try to calculate net torque about other points, maybe it's possible.\[T_2 \cos \phi (L) = mg(x)\]\[T_1 \cos \theta (L) = mg(L-x)\]Again, dunno if anything changes.

ganeshie8 (ganeshie8):

halliday http://as.wiley.com/WileyCDA/WileyTitle/productCd-EHEP002531.html

Parth (parthkohli):

Ah lol, I have a copy. Never opened the question bank.

ganeshie8 (ganeshie8):

it is a good book at the moment, i am reviewing Chapter 12 Equilibrium and Elasticity

ganeshie8 (ganeshie8):

I think we can express both T1 and T2 in terms of m, then plug it in the third eqn

Parth (parthkohli):

Haha, well, it works.\[mg =\left(\cos \phi + \cot \theta \sin \phi \right) T_2 \]

Parth (parthkohli):

\[\frac{mg}{T_2} = \frac{L\cos \phi }{x}\]

Parth (parthkohli):

\[x = \frac{L\cos \phi}{ \cos \phi + \cot \theta \sin \phi }\]